OpenStudy (anonymous):

a plane travels 228 miles in 1.8 hours against the wind. on the return trip, it travels the same 228 miles in 1 hour. Find the speed of the wind

OpenStudy (anonymous):

I would appreciate your help

OpenStudy (littlebird):

Do they give the wind speed?

OpenStudy (anonymous):

No they do not

OpenStudy (anonymous):

Let s and w be the plane's speed and the wind speed respectively. Solve the following for w: {18/10 (s - w) = 228, 1 (s + w) = 228} w = 152/3 mph

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

Your welcome.